OpenStudy (anonymous):

the math problem is 20/9*x over 4/3*3/2

OpenStudy (anonymous):

use equation to explain ur prob

OpenStudy (anonymous):

20/9 * X /4/3 * 3/2

OpenStudy (anonymous):

No instructions. Is this set to 0, solve for x?

OpenStudy (anonymous):

\[\frac{\frac{20}{9*x}}{\frac{4}{3}*\frac{3}{2}}=\frac{10}{9 x} \] Not sure if that is the intended expression.